OpenStudy (e.cociuba):

Solve the equation by completing the square: a^2 - 2a - 8= 0

jimthompson5910 (jim_thompson5910):

a^2 - 2a - 8 = 0 a^2 - 2a = 8 a^2 - 2a + 1 = 8 + 1 ... see note below (a - 1)^2 = 9 I'll let you finish up Note: on this line I took half of -2 to get -1, then squared it to get +1. I then added it to both sides.